7. The following table presents data on three leadin indicators for a three-month period. Construct the composite index (with each indicator assigned equal weight) and the diffusion index. Leading Leading Leading Month Indicator A Indicator B Indicator C 100 110 120 200 230 240 30 27 2
Answer #1

Solution:-

The composite index for the month from 2 to 3

We have the total percentage change of each series relative to the first month=40

Their average is 40/3=13.33

Now the composite index for the month of 3 is 13.33+100=113.33

Also the diffusionindex for the month 3

Procedure

The first step in computing the diffusion indexes is to calculate if a component increased, decreased, or had no change. Components that rise more than 0.05 percent are given a value of 1, components that change less than 0.05 percent are given a value of 0.5, and components that fall more than 0.05 percent are given a value of 0. Next, sum the values of the components. Third, divide by the number of components. Finally, multiply by 100.

In the month 3 all the changes are greater than 05 therefore their sum=1+1+1=3

Their average is 3/3 =1 In terms of percentage it is 1*100=100%

Therefore the diffusion index table and composite index table is

Month composite index diffusion index
1 100
2 105 66.7
3 113.33 100

Thanks